# HybridVLA: Collaborative Diffusion and Autoregression in a Unified Vision-Language-Action Model

### v2

@@ Abstract
Recent advancements in vision-language models (VLMs) for common-sense reasoning have led to the development of vision-language-action (VLA) models, enabling robots to perform generalized manipulation. Although existing autoregressive VLA methods leverage large-scale pretrained knowledge, they disrupt the continuity of actions. Meanwhile, some VLA methods incorporate an additional diffusion head to predict continuous actions, relying solely on VLM-extracted features, which limits their reasoning capabilities. In this paper, we introduce HybridVLA, a unified framework that seamlessly integrates the strengths of both autoregressive and diffusion policies within a single large language model, rather than simply connecting them. To bridge the generation gap, a collaborative training recipe is proposed that injects the diffusion modeling directly into the next-token prediction. With this recipe, we find that these two forms of action prediction not only reinforce each other but also exhibit varying performance across different tasks. Therefore, we design a collaborative action ensemble mechanism that adaptively fuses these two predictions, leading to more robust control. In experiments, HybridVLA outperforms previous state-of-the-art VLA methods across various simulation and real-world tasks, including both single-arm and dual-arm robots, while demonstrating stable manipulation in previously unseen configurations.
Figure 1: HybridVLA innovatively integrates diffusion and autoregressive action prediction within a single LLM, fully leveraging the continuity and probabilistic nature of diffusion alongside the reasoning capabilities of autoregressive modeling. It undergoes pretraining on large, diverse, cross-embodied real-world robotic datasets and is further fine-tuned on both simulation and self-collected real-world data. HybridVLA achieves remarkable performance across various tasks, demonstrating strong generalization to unseen manipulated objects, backgrounds, spatial positions, and lighting conditions.
@@ 1 Introduction
Recently, vision-language models (VLMs) [58, 3, 51, 112, 5, 21, 114, 113] have demonstrated exceptional abilities in instruction-following and common-sense reasoning, driven by pretraining on internet-scale image-text pairs. Building on this success, several studies have extended VLMs into vision-language-action (VLA) models, enabling them to generate action plans[1, 18, 34, 7] or predict SE(3) poses [10, 47, 56]. VLA models enable robots to interpret visual observations and language conditions, generating generalizable actions for control. Therefore, effectively harnessing the inherent capabilities of VLMs to develop VLA models for stable manipulation in dynamic environments is essential [55].
On the one hand, some existing VLA methods [10, 56, 47, 74] quantize continuous actions into discrete bins, replacing part of the vocabulary in large language models (LLMs). These autoregressive approaches mimic the next-token prediction of VLMs, effectively harnessing their large-scale pretrained knowledge while preserving reasoning capabilities. Although such methods enable generalized manipulation skills [47], the quantization process disrupts the continuity of action poses [97]. On the other hand, some VLA approaches [60, 33, 54, 100] incorporate a policy head (e.g., MLP or LSTM [25]) to transform LLM output embeddings into continuous action poses. However, these regressive methods overlook the scalability of the policy head and fail to incorporate probabilistic action representations [52, 14].
@@
Building on the success of diffusion models in content generation [31, 32, 73, 80, 28], diffusion policies have recently been introduced in robotic imitation learning [14, 110, 45, 71, 79, 104]. Unlike regressive deterministic policy heads, $\pi_{0}$[8], CogACT [52], and DiVLA [97] incorporate a diffusion head after VLMs, leveraging probabilistic noise-denoising for action prediction. While diffusion-based VLA methods enable precise manipulation, the diffusion head operates independently of the VLM, relying solely on VLM-extracted features as input conditions. Consequently, it fails to fully leverage the VLM’s reasoning capabilities. Given these advantages and limitations, a question arises: “How can we elegantly construct a unified VLA model that seamlessly integrates the strengths of both autoregressive and diffusion policies, rather than simply concatenating them?
@@
To achieve this, we propose HybridVLA, a unified framework that equips VLMs with both diffusion and autoregressive action prediction capabilities, enabling robust execution across diverse and complex manipulation tasks. Unlike previous diffusion-based VLA methods [8, 52], which append independent diffusion head after the LLM, HybridVLA seamlessly integrates diffusion modeling into the autoregressive next-token prediction within a single LLM, as shown in Figure 1. Specifically, we introduce a collaborative training recipe that encodes the diffusion-noised action as a continuous vector and projects it into the LLM’s word embedding space. To ensure consistency when combining the two generation methods, a token sequence formulation is designed to systematically organize multimodal input, diffusion action, and autoregressive action tokens while linking them through specialized marker tokens. With this design, we observe that the two action predictions not only reinforce each other but also exhibit varying performance across different tasks. For instance, the diffusion prediction excels in intricate tasks, while the autoregressive prediction performs better in tasks requiring rich semantic understanding. Therefore, a collaborative action ensemble mechanism is proposed, where the two predictions are adaptively fused based on autoregressive action token confidence, improving robustness in manipulation.
@@
To enhance HybridVLA’s generalization, we employ a step-by-step training approach [8], first performing large-scale pretraining, followed by fine-tuning on downstream tasks, as shown at the top of Figure 1. In addition to being initialized with parameters from a pretrained VLM [44], our model undergoes further pretraining on large, diverse, cross-embodiment robotic datasets, including Open X-Embodiment [69], DROID [46], and ROBOMIND [101], encompassing 760K trajectories and over 10K A800 GPU training hours. Subsequently, HybridVLA is fine-tuned on high-quality simulation data [36] and self-collected real-world data, achieving state-of-the-art (SOTA) manipulation performance across a variety of tasks with both single-arm and dual-arm robots. Meanwhile, HybridVLA demonstrates sufficient generalization capabilities to unseen manipulated objects, backgrounds, spatial positions, and lighting conditions during real-world testing, highlighting the effectiveness of our collaborative model design and training recipe. To optimize inference speed, we introduce HybridVLA-dif, which integrates diffusion and autoregressive generation during training but relies exclusively on diffusion-based actions for inference at 9.4 Hz. In summary, our contributions are as follows:
@@
  • We propose HybridVLA, a unified model that seamlessly integrates diffusion and autoregressive action prediction within a single LLM, effectively combining the continuous nature of diffusion-based actions with the reasoning capabilities of autoregressive generation.
@@
  • We introduce a collaborative training recipe that bridges the gap between the two action generation approaches, enabling mutual reinforcement. Additionally, we propose a collaborative action ensemble mechanism that adaptively fuses diffusion- and autoregressive-based actions, enhancing manipulation robustness.
@@
  • Our proposed method achieves SOTA performance across diverse tasks while demonstrating strong generalization to unseen configurations.
@@ 2 Related Work
Traditional robotic manipulation primarily relies on state-based reinforcement learning [4, 23, 42, 109], whereas recent approaches [92, 20, 9, 14] integrate visual observations for imitation learning. Building on the strong reasoning capabilities of vision-language models (VLMs) [58, 3, 51, 41, 27, 50], recent research has integrated them into robotic manipulation [53, 106, 107, 59].
@@
Vision-language-action (VLA) models. Some studies [1, 18, 34, 35] enable robots to interpret both language and visual observations, automatically generating task plans. Meanwhile, vision-language-action (VLA) models leverage the inherent reasoning abilities of VLMs to predict low-level SE(3) poses. Specifically, RT2[10] quantizes 7-DoF actions into discrete bins for autoregressive pose prediction. Building on this, ManipLLM[56] incorporates affordance priors through chain-of-thought reasoning, while OpenVLA[47] performs large-scale pretraining on the Open X-Embodiment dataset[69], enhancing generalization capabilities. FAST [74] applies the discrete cosine transform to enable fast and scalable training of autoregressive-based VLA models. To support continuous action prediction, some VLA approaches [60, 33, 54, 100] incorporate a policy head, such as an MLP or LSTM [25], and use regression loss for imitation learning. However, quantization in autoregressive methods disrupts action continuity, while regressive methods fail to incorporate probabilistic action representations.
@@
Diffusion models in robotics. Building on the success of diffusion models in content generation [31, 32, 73, 80], diffusion policies have been applied in robotics, including reinforcement learning [2, 96], imitation learning [14, 72, 75, 79, 104], grasping [88, 94, 102], and motion planning [38, 82]. Following this, 3D Diffusion Actor [45] and DP3[14] employ diffusion models to interpret point cloud data. Octo[91] and RDT-1B[61] enhance a transformer model with a diffusion head to predict flexible actions.
@@
Diffusion-based VLA models. To integrate diffusion with VLMs, $\pi_{0}$[8] adds a diffusion expert head that generates actions through flow matching, while TinyVLA [98] incorporates a simple diffusion head after the lightweight VLM. CogACT[52] and DiVLA [97] decouple reasoning and action prediction into the VLM and an injected diffusion head, respectively. However, in these methods, the diffusion head functions as a separate module, relying solely on LLM-extracted features for action conditions, limiting its use of the VLM’s pretrained knowledge and reasoning capabilities. Beyond robotics, in general multimodal scenarios, some works [22, 99, 103, 105] jointly address multimodal understanding and generation, while others [115, 12, 116] incorporate diffusion into an autoregressive-based transformer. Unlike previous approaches that combine image and language generation while preserving their respective quality, HybridVLA introduces a robotic-specific collaborative training recipe that seamlessly incorporates diffusion-based action generation into autoregressive action generation within a single LLM, effectively leveraging the strengths of both methods to mutually enhance each other.
@@ 3 HybridVLA Method
Overview. Existing diffusion-based VLA methods [8, 97, 52] append a separate diffusion head after the VLM, using VLM-extracted features as conditions for the diffusion process. However, these methods fail to fully exploit the VLM’s inherent reasoning capabilities derived from internet-scale pretraining. In contrast, HybridVLA equips a single LLM with both diffusion and autoregressive action generation capabilities. To construct HybridVLA, we first describe the model architecture in Section 3.1. Since simply merging the two generation methods could cause inconsistency, we introduce a collaborative training recipe in Section 3.2. To further enhance robustness, we propose a collaborative action ensemble mechanism in Section 3.3.
Problem Statement. At time t, each demonstration consists of image observations $o_{t}$, language description $l_{t}$, and the current robot state $r_{t}$. Our model $\pi$ aims to predict action $a$ to control the robot arms, which can be formulated as:
$\pi:(o_{t},l_{t},r_{t})\rightarrow a_{t+1}$
Following [47, 52], the action $a$ represents the end-effector pose, which uses 7-DOF and 14-DOF for single-arm and dual-arm control, respectively. Each 7-DOF action includes 3-DOF for relative translation offsets ($[\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DOF for rotation (Euler angles $\in\mathbb{R}^{3}$), and 1-DOF for the gripper state (open/closed $\in\mathbb{R}^{1}$). The ground truth (GT) and the model-predicted action are in SE(3), formulated as:
$a=[\Delta x,\Delta y,\Delta z,Roll,Pitch,Yaw,0/1]$
Figure 2: HybridVLA Framework. The input data, regardless of modality, is encoded and concatenated into our formatted token sequence. To integrate diffusion into the LLM, HybridVLA simultaneously projects the denoising timestep and noisy actions into the token sequence. The marker tokens, $<$BOD$>$ (beginning of diffusion) and $<$EOD$>$ (end of diffusion), are designed to bridge the two generation methods. By employing collaborative training to explicitly incorporate knowledge from both generation methods, these two action types reinforce each other and are adaptively ensembled to control the robot arms. For HybridVLA’s output, continuous actions are generated through iterative denoising, while discrete actions are produced autoregressively, all within the next-token prediction process.
@@ 3.1 HybridVLA Architecture
This section presents the architecture and workflow of HybridVLA, available in two model sizes, using 2.7B and 7B large language models (LLMs). Following [47], HybridVLA inherits a base architecture adapted from the Prismatic VLMs [44], leveraging its internet-scale pretrained parameters. We first introduce the two basic components—vision encoders and the LLM—as shown in Figure 2.
Vision encoders. HybridVLA leverages powerful vision encoder combinations, such as DINOv2 [70] and SigLIP [111], to capture rich semantic features $f_{d}\in\mathbb{R}^{B\times N_{v}\times 1024}$ and $f_{s}\in\mathbb{R}^{B\times N_{v}\times 1152}$. $B$ and $N$ represent batch size and token dimension, respectively. These features are concatenated along the channel dimension to form $f_{v}\in\mathbb{R}^{B\times N_{v}\times 2176}$, which is subsequently projected into the LLM’s word embedding via a projection layer. HybridVLA(2.7B) uses only the CLIP [77] model as its vision encoder. When processing multi-view images, a shared vision encoder extracts features, which are then concatenated along the token dimension.
LLM. HybridVLA adopts 7B LLAMA-2 [93] as LLM, responsible for multi-modal understanding and reasoning. Language prompts are encoded into embedding space $f_{l}\in\mathbb{R}^{B\times N_{l}\times 4096}$ using the pre-trained tokenizer, then concatenated with visual tokens and input into LLM. The other specially designed LLM inputs are presented in the next section, and the output tokens are processed in two ways. First, diffusion-based action ($a^{d}_{t+1}$) generation through a denoising process, where an MLP maps the tokens into the action space. Second, autoregressive-based action ($a^{ar}_{t+1}$) generation is performed through a detokenizer [47], which also computes the mean confidence ($c^{ar}_{t+1}$) of the action tokens, serving as a guiding factor for the collaborative action ensemble. For HybridVLA (2.7B), the workflow remains the same as that of HybridVLA (7B) but utilizes the 2.7B Phi-2 [39] as the LLM. In the next section, we introduce how to simultaneously equip a single LLM with diffusion and autoregressive action generation capabilities.

### v3

@@ Abstract
A fundamental objective of manipulation policy design is to endow robots to comprehend human instructions, reason about scene cues, and execute generalized actions in dynamic environments. Recent autoregressive vision-language-action (VLA) methods inherit common-sense reasoning capabilities from vision-language models (VLMs) for next action-token prediction. However, these methods quantize actions into discrete bins, which disrupts the continuity required for precise control. In contrast, existing diffusion-based VLA methods incorporate an additional diffusion head to predict continuous actions solely conditioned on feature representations extracted by the VLM, without fully leveraging the VLM’s pretrained reasoning capabilities through token-level generation. To address these limitations, we introduce HybridVLA, a unified framework that absorbs the continuous nature of diffusion-based actions and the contextual reasoning of autoregression within a single large language model. To mitigate interference between the two generation paradigms, we propose a collaborative training recipe that seamlessly incorporates diffusion denoising into the next-token prediction process. With this recipe, we find these two action prediction methods not only reinforce each other but also exhibit varying strength across different tasks. Therefore, we design a collaborative action ensemble mechanism that adaptively fuses both predictions, leading to more robust control. HybridVLA outperforms previous state-of-the-art VLA methods by 14% and 19% in mean success rate on simulation and real-world tasks, respectively, while demonstrating stable manipulation in unseen configurations.
@@ 1 Introduction
Developing human-like robots capable of performing manipulation tasks demands intelligent policies driess2023palm ; ahn2022can ; huang2023voxposer . In dynamic and unstructured real-world environments, such policies need to interpret human instructions and generalize across a wide range of complex tasks li2024towards . Recently, vision-language models (VLMs) liu2024visual ; alayrac2022flamingo ; li2023blip ; zhang2023llama have brought forth dramatic breakthroughs in instruction following and common-sense reasoning, driven by pretraining on internet-scale image-text pairs. Building on this success, several studies have extended VLMs into vision-language-action (VLA) models, enabling them to predict low-level action poses for robotic manipulation brohan2023rt ; kim2024openvla ; li2024manipllm . This paradigm outlines a promising roadmap for building foundation models to facilitate generalist robots.
Figure 1: (a) Unlike recent diffusion-based VLA methods bjorck2025gr00t ; black2024pi_0 ; li2024cogact that attach a separate diffusion head after VLMs, (b) HybridVLA innovatively integrates diffusion and autoregressive action prediction within a single LLM, fully leveraging the continuity of diffusion and the reasoning capabilities of autoregressive modeling. It is pretrained on large, diverse, cross-embodiment real-world robot datasets and further fine-tuned on downstream, self-collected data. HybridVLA achieves remarkable performance across various tasks involving both single-arm and dual-arm robots.
@@
On the one hand, autoregressive VLA methods brohan2023rt ; li2024manipllm ; kim2024openvla ; pertsch2025fast emulate the reasoning paradigm of VLMs for next token prediction, effectively leveraging their large-scale pretrained knowledge and reasoning capabilities. While such methods enable generalized manipulation skills kim2024openvla , they quantize continuous actions into discrete bins by adding new embeddings into the vocabulary in large language models (LLMs), which disrupts the continuity of action pose and hinders precise control wen2024diffusion . On the other hand, building on the success of diffusion models in content generation ho2020denoising ; ho2022video ; peebles2023scalable ; rombach2022high , diffusion policies have been introduced in robotic imitation learning chi2023diffusion ; ze20243d ; ke20243d ; pearceimitating ; reuss2023goal ; xian2023chaineddiffuser . Recent diffusion-based VLA methods black2024pi_0 ; li2024cogact ; wen2024diffusion ; bjorck2025gr00t incorporate a diffusion head after the VLM, leveraging probabilistic noise-denoising for action prediction. While these methods enable precise manipulation, the diffusion head operates independently of the VLM and lacks internet-scale pretraining. Moreover, since the head relies solely on VLM-extracted feature representations as input conditions, these methods fail to fully leverage the VLM’s pretrained reasoning capabilities through next-token prediction. Given these advantages and limitations, a question arises: “How can we elegantly construct a unified VLA model that seamlessly integrates the strengths of both autoregressive and diffusion policies, rather than simply concatenating them?"
@@
To this end, we propose HybridVLA, a unified framework that equips VLMs with both diffusion and autoregressive action prediction capabilities, enabling mutual reinforcement between them to facilitate robust execution across diverse tasks. As shown in Figure 1, unlike previous diffusion-based VLA methods black2024pi_0 ; li2024cogact that append an independent diffusion head after the LLM (Figure 1 (a)), we introduce a collaborative training recipe that seamlessly integrates diffusion denoising into the autoregressive next-token prediction process within a single LLM backbone (Figure 1 (b)). Specifically, since the token representations of discrete autoregressive tokens and continuous diffusion latents are inconsistent, a token sequence formulation is designed to systematically organize multimodal inputs, diffusion tokens, and autoregressive tokens, which are linked through specialized marker tokens. Under our proposed collaborative optimization, as both generation methods share the LLM backbone, HybridVLA explicitly captures the continuous action representations from diffusion modeling along with the pretrained semantic reasoning of autoregressive generation, allowing the two paradigms to reinforce each other. Meanwhile, we observe that diffusion generation excels in intricate tasks, while autoregression performs better in tasks requiring rich semantic understanding. Therefore, a collaborative action ensemble mechanism is proposed, where the two predictions are adaptively fused based on autoregressive action token confidence, improving robustness in manipulation.
@@
To enhance generalization capability, we initialize HybridVLA with an internet-scale pretrained VLM karamcheti2024prismatic , and design a step-by-step training approach black2024pi_0 ; kim2024openvla . As shown in Figure 1, our model undergoes further pretraining on large, diverse, cross-embodiment robotic datasets, including Open X-Embodiment o2023open , DROID khazatsky2024droid , and ROBOMIND wu2024robomind , covering 760K trajectories and over 10K A800 GPU training hours. Subsequently, HybridVLA is fine-tuned on self-collected simulation data james2020rlbench and real-world data, achieving state-of-the-art (SOTA) manipulation performance across a variety of tasks with both single-arm and dual-arm robots. Meanwhile, HybridVLA demonstrates sufficient generalization capabilities to unseen manipulated objects, backgrounds, spatial positions, and lighting conditions during real-world testing, highlighting the effectiveness of our collaborative model design and training recipe. To optimize inference speed, we also introduce the HybridVLA-dif (7B) variant, which integrates diffusion and autoregressive generation during training but relies exclusively on diffusion-based actions for inference at 9.4 Hz. Our contributions are as follows:
@@
  • We propose HybridVLA, a unified model that seamlessly integrates diffusion and autoregressive action generation within a single LLM, effectively absorbing the continuous nature of diffusion-based actions and the contextual reasoning of autoregressive generation, thereby enabling mutual reinforcement and improving manipulation robustness.
@@
  • We introduce a collaborative training recipe that bridges the gap between the two action generation approaches, enabling mutual reinforcement through a shared LLM backbone. Additionally, we propose a collaborative action ensemble mechanism that adaptively fuses diffusion- and autoregressive-based actions, enhancing manipulation robustness.
@@
  • Our proposed method achieves SOTA performance across diverse tasks while demonstrating strong generalization to several unseen configurations.
@@ 2 Related Work
Traditional robotic manipulation primarily relies on state-based reinforcement learning andrychowicz2020learning ; geng2022end ; joshi2020robotic ; yarats2021mastering , whereas recent approaches torabi2018behavioral ; fu2024mobile ; brohan2022rt ; chi2023diffusion integrate visual observations for imitation learning. Building on the strong reasoning capabilities of vision-language models (VLMs) liu2024visual ; alayrac2022flamingo ; li2023blip ; li2024llava , recent research has integrated them into robotic manipulation licrayonrobo ; xiong2024autonomous ; xu2024naturalvlm ; liu2024self .
@@
Vision-language-action (VLA) models. Some studies ahn2022can ; driess2023palm ; huang2023voxposer ; huang2024rekep enable robots to interpret both language and visual observations, automatically generating task plans. Meanwhile, vision-language-action (VLA) models leverage the inherent reasoning abilities of VLMs to predict low-level SE(3) poses. Specifically, RT2brohan2023rt quantizes 7-DoF actions into discrete bins for autoregressive pose prediction. Building on this, ManipLLMli2024manipllm incorporates affordance priors through chain-of-thought reasoning, while OpenVLAkim2024openvla performs large-scale pretraining on the Open X-Embodiment dataseto2023open . FAST pertsch2025fast applies the discrete cosine transform to enable fast and scalable training of autoregressive-based VLA models. To support continuous action prediction, some VLA approaches liu2024robomamba ; huang2024manipvqa ; li2023vision ; wu2023unleashing incorporate a policy head, such as an MLP or LSTM graves2012long , and use regression loss for imitation learning. However, quantization in autoregressive methods disrupts action continuity, while regressive methods fail to incorporate probabilistic action representations.
@@
Diffusion models in robotics. Building on the success of diffusion models in content generation ho2020denoising ; ho2022video ; peebles2023scalable ; rombach2022high , diffusion policies have been applied in robotics, including reinforcement learning ajay2022conditional ; wang2022diffusion , imitation learning chi2023diffusion ; pearce2023imitating ; prasad2024consistency ; reuss2023goal ; xian2023chaineddiffuser , grasping simeonov2023shelving ; urain2023se ; wu2023learning , and motion planning janner2022planning ; saha2024edmp . Following this, 3D Diffusion Actor ke20243d and DP3 chi2023diffusion employ diffusion models to interpret point cloud data. Octo team2024octo and RDT-1B liu2024rdt augment a transformer with a diffusion head to predict flexible actions.
@@
Diffusion-based VLA models. To integrate diffusion with VLMs, $\pi_{0}$ black2024pi_0 adds a diffusion expert head that generates actions through flow matching, while TinyVLA wen2024tinyvla incorporates a simple diffusion head after the lightweight VLM. CogACT li2024cogact and DiVLA wen2024diffusion decouple reasoning and action prediction into the VLM and an injected diffusion head, respectively. Following this architecture, some works bjorck2025gr00t ; bu2025agibot ; figure2024helix introduce a dual-system design to enable control at different frequencies. However, in these methods, the diffusion head operates as a separate module and treats the VLM as a multimodal feature extractor, limiting its ability to fully exploit pretrained reasoning capabilities through next-token prediction. In general scenarios, some works ge2024seed ; wu2024janus ; wu2024vila ; xie2024show jointly tackle multimodal understanding and generation, while others zhao2024monoformer ; chen2024diffusion ; zhou2024transfusion integrate diffusion into autoregressive transformers. Unlike prior methods focused on image and language generation quality, HybridVLA introduces a robotics-specific collaborative training strategy that integrates diffusion action generation into next-token prediction within a single LLM, enabling mutual enhancement.
@@ 3 HybridVLA Method
Overview. Existing diffusion-based VLA methods black2024pi_0 ; wen2024diffusion ; li2024cogact append a separate diffusion head after the VLM. However, these methods overlook the LLM’s core contextual reasoning mechanism (next-token prediction) acquired through internet-scale pretraining, since the head relies solely on VLM-extracted multimodal features from a single forward pass as diffusion conditions. In contrast, HybridVLA injects diffusion denoising into the next-token prediction process, equipping a single LLM with both diffusion and autoregressive action generation capabilities. To construct HybridVLA, we first describe the model architecture in Section 3.1. Since simply merging the two generation methods could cause inconsistency, we introduce a collaborative training recipe in Section 3.2. To further enhance robustness, we propose a collaborative action ensemble mechanism in Section 3.3.
Problem Statement. At time t, each demonstration consists of image observations $o_{t}$, language description $l_{t}$, and the current robot state $r_{t}$. Our model $\pi$ aims to predict action $a$ to control the robot arms, which can be formulated as: $\pi:(o_{t},l_{t},r_{t})\rightarrow a_{t+1}$. Following kim2024openvla ; li2024cogact , the action $a$ represents the end-effector pose, which uses 7-DOF and 14-DOF for single-arm and dual-arm control, respectively. Each 7-DOF action includes 3-DOF for relative translation offsets ($[\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DOF for rotation (Euler angles $\in\mathbb{R}^{3}$), and 1-DOF for the gripper state (open/closed $\in\mathbb{R}^{1}$). The ground truth (GT) and the model-predicted action are in SE(3), formulated as: $a=[\Delta x,\Delta y,\Delta z,Roll,Pitch,Yaw,0/1]$.
Figure 2: HybridVLA Framework. All multimodal inputs are encoded into tokens and subsequently organized into our designed token sequence formulation within the LLM’s embedding space. For diffusion tokens, HybridVLA simultaneously projects the denoising timestep and noise into continuous vector representations. During inference, we adopt DDIM song2020denoising with four sampling steps, where the corresponding noisy samples are iteratively fed into the LLM to predict the noise at each step. The marker tokens, $<$BOD$>$ (Beginning of Diffusion) and $<$EOD$>$ (End of Diffusion), are introduced to bridge the two generation paradigms. Subsequently, autoregressive actions are generated via standard next action-token prediction, explicitly conditioned on the preceding tokens. Our collaborative training recipe integrates knowledge from both generation paradigms into the shared LLM, enabling them to reinforce each other and be adaptively ensembled for robot arm control.
@@ 3.1 HybridVLA Architecture
Pretrained VLM base. This section presents the architecture and workflow of HybridVLA, available in two model sizes, using 7B and 2.7B large language models (LLMs). Following kim2024openvla , both HybridVLA(7B) and HybridVLA(2.7B) inherit the base architecture from Prismatic VLMs karamcheti2024prismatic , initializing with the corresponding large-scale pretrained VLM parameters. We first introduce the two basic components, vision encoders and the LLM, as shown in Figure 2.
Vision encoders. HybridVLA leverages powerful vision encoder combinations, such as DINOv2 oquab2023dinov2 and SigLIP zhai2023sigmoid , to capture rich semantic features $f_{d}\in\mathbb{R}^{B\times N_{v}\times 1024}$ and $f_{s}\in\mathbb{R}^{B\times N_{v}\times 1152}$. $B$ and $N$ represent batch size and token sequence length, respectively. These features are concatenated along the channel dimension to form $f_{v}\in\mathbb{R}^{B\times N_{v}\times 2176}$, which is subsequently projected into the LLM’s word embedding via a projection layer. HybridVLA(2.7B) uses only the CLIP radford2021learning model as its vision encoder. When processing multi-view images, a shared vision encoder extracts features, which are then concatenated along the token dimension.
LLM. HybridVLA adopts 7B LLAMA-2 touvron2023llama as LLM, responsible for multimodal understanding and reasoning. Language prompts are encoded into embedding space $f_{l}\in\mathbb{R}^{B\times N_{l}\times 4096}$ using the pre-trained tokenizer, then concatenated with visual tokens and input into LLM. The other specially designed LLM inputs (e.g., diffusion noise) are presented in the next section, and the output tokens are processed in two ways. First, diffusion-based action ($a^{d}_{t+1}$) generation through a denoising process, where an MLP maps the tokens into the action space. Second, autoregressive-based action generation ($a^{ar}_{t+1}$) is performed using a detokenizer kim2024openvla , which also computes the mean confidence ($c^{ar}_{t+1}$) of the predicted tokens, serving as a guiding factor for the collaborative action ensemble. For HybridVLA (2.7B), the workflow remains the same as that of HybridVLA (7B) but utilizes the 2.7B Phi-2 javaheripi2023phi as the LLM. In the next section, we introduce how to simultaneously equip a single LLM with diffusion and autoregressive action generation capabilities.